THE ZÜRICH HIGH RESOLUTION COLLISIONAL-COOLING CELL-FTIR SETUP: ROVIBRATIONAL SPECTROSCOPY OF METHANOID MOLECULES BETWEEN 60 AND 300 K

\underline{SIEGHARD ALBERT}, KAREN KEPPLER ALBERT, SIGURD BAUERECKER AND MARTIN QUACK, Physical Chemistry, ETH Zürich, CH-8093 Zürich, Switzerland.

A complete understanding of the absorption behavior of the Earth’s atmosphere and of the atmospheres of the giant planets (Saturn and Jupiter) and their moons (Titan) requires a detailed spectroscopic investigation of methanoid molecules over full atmospheric temperature ranges. For that reason we have interfaced a collisional and enclosive cooling cell^{a} based on White-type multireflection optics to our FTIR IFS125 HR prototype 2001 spectrometer^{b} (MOPD=10 m). This cell makes it possible to record spectra at high spectral resolution in the temperature range 4-400 K with absorption path lengths up to 20 m. We have recorded the spectra of CH_{4}^{c} and its isotopomers^{d} and of CHF_{3}^{e}, CHClF_{2}^{f} and CHCl_{2}F^{a} between 2000 and 6000 cm^{−1} in the temperature range 60-300 K. The spectra were recorded at resolutions ranging from 0.0015 to 0.004 cm^{−1}. We will present an analysis of the spectra of CHCl_{2}F in the 2ν_{3} and 3ν_{3} regions and an initial assignment of the resonance system 2ν_{3}/ν_{3} + ν_{8}/2ν_{8} of CHClF_{2}. We discuss coincidences with CO_{2} laser lines in the 2ν_{3} region of the isotopic chiral molecule CH^{35}Cl^{37}ClF. Doppler-free quasi-resonantly enhanced ultra-high two-photon absorption experiments may be carried out in this region to study parity violation^{g} in this molecule. In addition, we show CO nanoparticles recorded at 6 K ^{h}.

aS. Albert, S. Bauerecker, M. Quack and A. Steinlin, Mol. Phys. in press.

bS. Albert, K.K. Albert, and M. Quack, Trends in Optics and Photonics 84, 177 (2003).

cA. Nikitin, V. Boudon, J.-P. Champion, S. Albert, S. Bauerecker, M. Quack and L. R. Brown, 61st International Symposium on Molecular Spectroscopy, Columbus, Ohio, USA, June 19-23, 2006, Talk RC11.

dO.N. Ulenikov, E.S. Bekthereva, S. Albert, S. Bauerecker, H. Hollenstein and M. Quack, to be published.

eS. Albert, S. Bauerecker and M. Quack, 18th Colloq. on High Res. Mol. Spectrosc., Dijon, 8-12 September 2003, page 352.

fS. Albert, H. Hollenstein, M. Quack and M. Willeke, Mol. Phys. 102, 1671 (2004).

gM. Quack, Angew. Chem. Intl. Ed. Engl. 114, 4812 (2002), M. Quack and J. Stohner, J. Chem. Phys. 119, 11228 (2003), R. Berger, G. Laubender, M. Quack, A. Sieben, J. Stohner and M. Willeke, Angew. Chem. Int. Ed. Eng. 44, 3623 (2005).

hS. Bauerecker, M. Taraschewski, C. Weitkamp and H. Cammenga, Rev. Sci. Instrum. 72, 3946 (2001).